OpenStudy (anonymous):

A math teacher gave her class two tests. Twenty-five percent of the class passed both tests and 42 percent of the class passed the first test. What percent of those who passed the first test also passed the second test? Answer 5% 10.5% 15.5% 25% Question 49 The probability that it is Friday and that a student is absent is 0.03. Since there are five school days in a week, the probability that it is Friday is 0.2. What is the probability that a student is absent given that today is Friday? Answer 0.10 0.15 0.20 0.25

OpenStudy (anonymous):

i think the answer to 1 is (25%) but im not sure about 2

OpenStudy (anonymous):

@Mariam.A

OpenStudy (anonymous):

Can someone just help me plzzzzzzzzzzz

OpenStudy (kropot72):

Question 49 Let F be the event 'it is Friday' and let A be the event 'student is absent'. \[\large P(F \cap A)=0.03\] It is given that \[\large P(F)=0.2\] Therefore \[\large P(A)=\frac{P(F \cap A)}{P(F)}=\frac{0.03}{0.2}=you\ can\ calculate\]

OpenStudy (anonymous):

is the answer 0.15?

OpenStudy (kropot72):

Yes, you are correct.

OpenStudy (anonymous):

yes kropot72 i have a few more Q's can u help me answer them too?

OpenStudy (anonymous):

for the second on use conditional probability rule P(A/F)=P(A and F)/P(A) =0.03/0.2 =0.15

OpenStudy (kropot72):

@MeggyMeg Sorry, I must log out now :(

OpenStudy (anonymous):

oh its okay :]

OpenStudy (anonymous):

im gonna give u a medal

OpenStudy (anonymous):

Mariam can u helo me?

OpenStudy (anonymous):

yeah I'll try m not a probability genius but I think I can manage :)

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

What is the mode of this set of data: 10, 11, 11, 12, 12, 12, 15, 15, 19, 20, 21, 22, 22? Answer 10 12 13 15 1 points Question 37 A football player ran the ball for 40 yards, 45 yards, 48 yards, 52 yards, and 105 yards. What is the outlier? Answer 105 52 40 65 1 points Question 38 A student had the following scores on quizzes: 26, 28, 34, and 37. What score would be needed on one additional quiz for this student to have a mean score of 31? Answer 30 31 34 38

OpenStudy (anonymous):

u dont have to answer question 1 i know that one

OpenStudy (anonymous):

ok for the second one the outlier is usually the number out of place or way out of the rest of the group, so from this definition which one do you think it should be?

OpenStudy (anonymous):

is it a 105?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

for the last one add up all the score and make the unknown score x, then divide this by the number of score equating it to 31 to make a formula like this\[\frac{ 26+28+34+37+x }{ 5 }=31\]

OpenStudy (anonymous):

try solve this and let me know how you go

OpenStudy (anonymous):

I mean solve for x which would be the additional score needed